# Bohr–Sommerfeld rules for systems

### v1

@@ 2. Microlocal preparation
Let $\mu_{1}=\mu_{1}^{\prime}+\mu_{1}^{\prime\prime}+\mu_{1}^{\prime\prime\prime}$ be as in (10). Let $H$ be modified using Lemmas 2.1 and 2.2 if necessary. Then $\operatorname{Spec}(H^{w}(x,hD))\cap[E_{-},E_{+}]$ is described by the Bohr–Sommerfeld rule (2), where $f_{1}$ in (4) is given by
@@ 3. The subprincipal term
Here we describe $\mu_{1}=\mu_{1}^{\prime}+\mu_{1}^{\prime\prime}+\mu_{1}^{\prime\prime\prime}$ in terms of spherical coordinates $\theta$ and $\phi$ in (11). To the eigenvalues $\lambda_{\pm}$ we choose corresponding eigenvectors $u_{\pm}$ of $H_{0}$ in (12), which are independent of $p_{0}$ and well-defined when $\lVert P\rVert\neq 0$, such that
@@
Lemma 3.3.
[…]
Proof.
Write $e=u_{\pm}$. By the chain rule we have
$\langle e_{x}^{\prime},e_{\xi}^{\prime}\rangle=\theta_{x}^{\prime}\theta_{\xi}^{\prime}\lVert e_{\theta}^{\prime}\rVert^{2}+\phi_{x}^{\prime}\phi_{\xi}^{\prime}\lVert e_{\phi}^{\prime}\rVert^{2}+\phi_{x}^{\prime}\theta_{\xi}^{\prime}\langle e_{\phi}^{\prime},e_{\theta}^{\prime}\rangle+\phi_{\xi}^{\prime}\theta_{x}^{\prime}\langle e_{\theta}^{\prime},e_{\phi}^{\prime}\rangle.$
Since $\operatorname{Im}\langle e_{\theta}^{\prime},e_{\phi}^{\prime}\rangle=-\operatorname{Im}\langle e_{\phi}^{\prime},e_{\theta}^{\prime}\rangle$ we get
[…]
A straightforward calculation using (18) shows that $\operatorname{Im}\langle e_{\phi}^{\prime},e_{\theta}^{\prime}\rangle=\pm\frac{1}{4}\sin(\theta)$ when $e=u_{\pm}$. The result now follows by inserting the expression for $\lambda_{\pm}$. ∎
@@
From the proof of Lemma 3.3 we have $\operatorname{Im}\langle e_{x}^{\prime},e_{\xi}^{\prime}\rangle=\pm\frac{1}{4}\sin(\theta)\{\theta,\phi\}$ which shows that this expression for the Berry curvature is the same as (15).
@@
Remark 3.6 (Rammal-Wilkinson’s phase).
@@
As we have seen, the right-hand side is the Berry curvature scalar density in phase space coordinates (cf. (21)). Let’s denote it by $f_{\pm}$, then, by Lemmas 3.1 and 3.3 we have $\mu_{1}^{\prime}=\pm\lVert P\rVert f_{\pm}$ and $\mu_{1}^{\prime\prime\prime}=\frac{1}{2}(p_{0}\pm\lVert P\rVert)f_{\pm}$, which we summarize as
@@
Similar to the case of the Berry phase, we get the following rigidity result for the Rammal-Wilkinson phase.
@@
Let $v_{+},w_{+}$ be any smooth linearly dependent normalized vectors with orthogonal linearly dependent normalized vectors $v_{-},w_{-}$ in a neighbourhood of a simple closed smooth curve $\gamma$, then their Rammal-Wilkinson phases computed from (22) coincide.
@@
$\mu_{1}=\lVert P\rVert\frac{\sin(\theta)}{2}\{\theta,\phi\}\pm\frac{1-\cos(\theta)}{2}\{\lambda_{\pm},\phi\}+(\pm p_{0}+\lVert P\rVert)\frac{\sin(\theta)}{4}\{\theta,\phi\}.$ (23)
Combining the first and third term, the stated formulas for $\theta_{B}$ and $\theta_{\mathit{RW}}$ follow in view of Remarks 3.4 and 3.6.
@@
When integrating the Berry connection or the Berry curvature scalar density of these eigenvectors over an energy level curve $E>0$, then for $\varepsilon>0$ small enough, the corresponding smooth eigenvector can be replaced by a smooth section of eigenvectors on that energy level curve of the unperturbed system. More precisely, there is the smooth choice of eigenvectors of the perturbed system $v_{\varepsilon}$ and the smooth choice of the unperturbed system $v$, i.e., by $u_{\pm}$ from (18). On an energy level set $\gamma$ they agree up to a continuous phase factor $e^{i\eta}:\gamma\to\mathbb{S}^{1},$ the complex unit circle. It follows from Lemmas 3.5 and 3.7 that the two computations agree up to a term $2\pi n$ for $n\in\mathbb{Z}$ which does not affect the Bohr–Sommerfeld rule. This also shows the invariance statement in Theorem 1.2 and the proof is complete. ∎
@@ 4. Quantized geometric phase corrections
$q_{1}$ $=p_{1}\bigg{(}c_{3}+\frac{c_{2}^{2}(1-c_{3})}{c_{1}^{2}+c_{2}^{2}}\bigg{)}-p_{2}\frac{c_{1}c_{2}(1-c_{3})}{c_{1}^{2}+c_{2}^{2}}-c_{1}p_{3},$ (24)
@@
where $\theta_{B}$ is the Berry phase. In particular, with $\mu_{1}=\mu_{1}^{\prime}+\mu_{1}^{\prime\prime}+\mu_{1}^{\prime\prime\prime}$ as in (10) we have $\mu_{1}^{\prime}=\mu_{1}^{\prime\prime\prime}=0$.
We note that Theorem 1.3 from the introduction is just a special case of Theorem 4.4. From Theorem 4.4 we also see that $H_{0}$ will always have trivial Berry curvature (thus leading to trivial Rammal-Wilkinson phase) if the components of $P=(p_{1},p_{2},p_{3})$ are linearly dependent over $\mathbb{R}$. To find an example of non-trivial Rammal-Wilkinson phase we must therefore look for a system where $P$ has linearly independent components. A simple example having both non-trivial Rammal-Wilkinson phase and non-trivial Berry phase is provided in §5.2.
@@ 5.1. The Jackiw-Rebbi model
$Q(x,hD_{x})\varphi=0\Rightarrow h\partial_{x}\varphi+m(x)\varphi=0.$
This shows that $\varphi(x)\propto\operatorname{exp}\left(-\int_{0}^{x}\frac{m(\xi)}{h}\ d\xi\right)\in L^{2}(\mathbb{R}),$ while the adjoint operator does not admit a $L^{2}$ zero mode. Thus, the Fredholm index is
@@
5.2. Non-trivial Rammal-Wilkinson phase
@@
From (23) we see that the terms $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime\prime}$ combine into
$\mu_{1}^{\prime}+\mu_{1}^{\prime\prime\prime}=\frac{3}{4}\lVert P\rVert\sin(\theta)\{\theta,\phi\}=\frac{3x^{2}}{4(x^{2}+\xi^{2}+x^{4})}.$
@@
$\int_{\gamma}(\mu_{1}^{\prime}+\mu_{1}^{\prime\prime\prime})\,dt=\frac{3}{4E^{2}}\int_{\{x^{2}+\xi^{2}+x^{4}=E^{2}\}}x^{2}\,dt.$
@@
$\int_{\gamma}(\mu_{1}^{\prime}+\mu_{1}^{\prime\prime\prime})\,dt=\frac{3}{E}\int_{0}^{x_{0}}\frac{x^{2}}{\sqrt{E^{2}-x^{2}-x^{4}}}\,dx$ (25)
@@
The Rammal-Wilkinson phase (25) is clearly non-zero, and can be computed numerically as a function of $E$. It is not quantized but depends continuously on $E$.
@@
Figure 3. Left: Comparison between spectral computation of the 14 smallest positive eigenvalues for $h=0.01$ compared with semiclassical approximation by $S_{0}$ and $S_{0}+hS_{1}$, respectively. The addition of the $S_{1}$ term leads to improvement away from zero energy. Right: We plot the non-quantized Berry and Rammal-Wilkinson phases.

### v2

@@ 2. Microlocal preparation
Let $\mu_{1}=\mu_{1}^{\prime}+\mu_{1}^{\prime\prime}$ be as in (10). Let $H$ be modified using Lemmas 2.1 and 2.2 if necessary. Then $\operatorname{Spec}(H^{w}(x,hD))\cap[E_{-},E_{+}]$ is described by the Bohr–Sommerfeld rule (2), where $f_{1}$ in (4) is given by
@@ 3. The subprincipal term
Here we describe $\mu_{1}=\mu_{1}^{\prime}+\mu_{1}^{\prime\prime}$ in terms of spherical coordinates $\theta$ and $\phi$ in (11). To the eigenvalues $\lambda_{\pm}$ we choose corresponding eigenvectors $u_{\pm}$ of $H_{0}$ in (12), which are independent of $p_{0}$ and well-defined when $\lVert P\rVert\neq 0$, such that
@@
Remark 3.5 (Rammal–Wilkinson’s phase).
@@
As we have seen, the right-hand side is the Berry curvature scalar density in phase space coordinates (cf. (20)). Let’s denote it by $f_{\pm}$, then $\mu_{1}^{\prime}=\pm\lVert P\rVert f_{\pm}$ by Lemma 3.1, so
@@
Similar to the case of the Berry phase, we get the following rigidity result for the Rammal–Wilkinson phase.
@@
Let $v_{+},w_{+}$ be any smooth linearly dependent normalized vectors with orthogonal linearly dependent normalized vectors $v_{-},w_{-}$ in a neighbourhood of a simple closed smooth curve $\gamma$, then their Rammal–Wilkinson phases computed from (21) coincide.
@@
$\mu_{1}=\lVert P\rVert\frac{\sin(\theta)}{2}\{\theta,\phi\}\pm\frac{1-\cos(\theta)}{2}\{\lambda_{\pm},\phi\}.$ (22)
The stated formulas for $\theta_{B}$ and $\theta_{\mathit{RW}}$ now follow in view of Remarks 3.3 and 3.5.
@@
When integrating the Berry connection or the Berry curvature scalar density of these eigenvectors over an energy level curve $E>0$, then for $\varepsilon>0$ small enough, the corresponding smooth eigenvector can be replaced by a smooth section of eigenvectors on that energy level curve of the unperturbed system. More precisely, there is the smooth choice of eigenvectors of the perturbed system $v_{\varepsilon}$ and the smooth choice of the unperturbed system $v$, i.e., by $u_{\pm}$ from (17). On an energy level set $\gamma$ they agree up to a continuous phase factor $e^{i\eta}:\gamma\to\mathbb{S}^{1},$ the complex unit circle. It follows from Lemmas 3.4 and 3.6 that the two computations agree up to a term $2\pi n$ for $n\in\mathbb{Z}$ which does not affect the Bohr–Sommerfeld rule. This also shows the invariance statement in Theorem 1.2 and the proof is complete. ∎
@@ 4. Quantized geometric phase corrections
$q_{1}$ $=p_{1}\bigg(c_{3}+\frac{c_{2}^{2}(1-c_{3})}{c_{1}^{2}+c_{2}^{2}}\bigg)-p_{2}\frac{c_{1}c_{2}(1-c_{3})}{c_{1}^{2}+c_{2}^{2}}-c_{1}p_{3},$
@@
By a simple calculation, this simplifies to
$q_{1}=p_{1}-\frac{c_{1}}{1+c_{3}}p_{3},\qquad q_{2}=p_{2}-\frac{c_{2}}{1+c_{3}}p_{3}.$ (23)
@@
where $\theta_{B}$ is the Berry phase. In particular, with $\mu_{1}=\mu_{1}^{\prime}+\mu_{1}^{\prime\prime}$ as in (10) we have $\mu_{1}^{\prime}=0$.
We note that Theorem 1.3 from the introduction is just a special case of Theorem 4.4. From Theorem 4.4 we also see that $H_{0}$ will always have trivial Berry curvature (thus leading to trivial Rammal–Wilkinson phase) if the components of $P=(p_{1},p_{2},p_{3})$ are linearly dependent over $\mathbb{R}$. To find an example of non-trivial Rammal–Wilkinson phase we must therefore look for a system where $P$ has linearly independent components. A simple example having both non-trivial Rammal–Wilkinson phase and non-trivial Berry phase is provided in §5.2.
@@ 5.1. The Jackiw-Rebbi model
$Q(x,hD_{x})\varphi=0\quad\Longrightarrow\quad h\partial_{x}\varphi+m(x)\varphi=0.$
This shows that $\varphi(x)\propto\operatorname{exp}\left(-\int_{0}^{x}\frac{m(\xi)}{h}\ d\xi\right)\in L^{2}(\mathbb{R}),$ while the adjoint operator does not admit an $L^{2}$ zero mode. Thus, the Fredholm index is
@@
5.2. Non-trivial Rammal–Wilkinson phase
@@
From (22) we see that
$\mu_{1}^{\prime}=\frac{1}{2}\lVert P\rVert\sin(\theta)\{\theta,\phi\}=\frac{x^{2}}{2(x^{2}+\xi^{2}+x^{4})}.$
@@
$\int_{\gamma}\mu_{1}^{\prime}\,dt=\frac{1}{2E^{2}}\int_{\{x^{2}+\xi^{2}+x^{4}=E^{2}\}}x^{2}\,dt.$
@@
$\int_{\gamma}\mu_{1}^{\prime}\,dt=\frac{2}{E}\int_{0}^{x_{0}}\frac{x^{2}}{\sqrt{E^{2}-x^{2}-x^{4}}}\,dx$ (24)
@@
The Rammal–Wilkinson phase (24) is clearly non-zero, and can be computed numerically as a function of $E$. It is not quantized but depends continuously on $E$.
@@
Figure 3. Left: Comparison between spectral computation of the 11 smallest positive eigenvalues for $h=0.01$ compared with semiclassical approximation by $S_{0}$ and $S_{0}+hS_{1}$, respectively. The addition of the $S_{1}$ term leads to improvement away from zero energy. Right: We plot the non-quantized Berry and Rammal–Wilkinson phases.